OpenStudy (anonymous):

4y 8 add and write in simplest form --- + --- y+2 y+2

OpenStudy (anonymous):

the denominators are same , so simply add , then take something common from numerator and try to cancel with denominator

OpenStudy (anonymous):

You lucky dog. The denominators are already the same! Because they are the same you can add the two together just like normal. But after that you still need to simplify by factoring the new top.

OpenStudy (anonymous):

i think im wrong but is it 2y +4 --------- y

OpenStudy (anonymous):

\[\frac{4y}{y+2}+\frac{8}{y+2}=\frac{4y+8}{y+2}\]

OpenStudy (anonymous):

ohhhh that makes sense! wow now i feel dumb